# The Automorphism Equivariant Hitchin Index

### v1

@@ 2 The Fixed Locus
Let $\Sigma$ be a compact Riemann surface with a finite subset $\delta$ of marked points. Our conventions and notation on quasi-parabolic Higgs bundles and parabolic Higgs bundles on $(\Sigma,\delta)$ agree with those in [46], except we only consider full flags and specialize to the rank two case. Notice that the real numbers $\{\alpha_{j}(y)\}_{j=1,2,y\in\delta}$ introduced in Section $1$ in [46] satisfy that $\alpha_{j}(y)=w_{j}(y),j=1,2$ for all marked points $y\in\delta$.
@@
Let $(E,\Phi)$ be a Higgs bundle with isomorphism class $\operatorname{\mathcal{E}}\in\mathcal{M}^{f}$. There is a unique $\langle f\rangle$-equivariant structure with generator $f_{E}:(E,\Phi)\rightarrow f^{*}(E,\Phi)$ covering the action of $f$ on $X$, such that $\det(E)$ is $\langle f\rangle$-equivariantly isomorphic to $\Lambda.$ Denote by $D_{E}$ the subset of $x\in X^{f}$ such that the eigenvalues of $f_{E}(x)$ are disctint, say $\sqrt{\alpha_{x}}\zeta^{i_{x}}$ and $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$ with $i_{x}\in\{1,...,(p-1)/2\}$. For each $x\in D_{E}$ let $F_{x}$ be the $\sqrt{\alpha_{x}}\zeta^{-i_{x}}$-eigenspace of $f_{E}(x)$. Define the set of flags $F=\{F_{x}\}_{x\in D_{E}}$. The isomorphism class of the $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E,F,\Phi,f_{E})$ on $(X,D_{E})$ depends only on $\operatorname{\mathcal{E}}$.
@@
where $f^{c}_{E}:\operatorname{End}_{0}(E)\rightarrow\operatorname{End}_{0}(f^{*}(E))$ is the isomorphism given by conjugation with $f_{E}$ and where $\tilde{f}^{*}(\Phi)$ is defined by equation (18). Notice that equation (19) is invariant under $f_{E}\mapsto\alpha\cdot f_{E}$ for any $\alpha\in\mathbbm{C}^{\times}$. Recall that since $(E,\Phi)$ is stable we have $f_{E}^{p}\in\operatorname{Aut}(E,\Phi)\simeq\mathbbm{C}^{\times}$. Thus, we can normalize to ensure $f_{E}^{p}=\operatorname{Id}$ and because of conjugation invariance, the equation (19) remain true. Thus $f^{*}([E,\Phi])=[(E,\Phi)]$ implies that there exists a lift $f_{E}$ of $f$ to $(E,\Phi)$. This is unique up to multiplying by an element of $\mu_{p}$. Now let $h:\det(E)\rightarrow\Lambda$ be an isomorphims (this exists by assumption). Let $f_{\Lambda}$ be the canonical $\langle f\rangle$-equivariant structure on $\Lambda$. Then $f_{1}:=h\circ\det(f_{E})\circ h^{-1}$ and $f_{\Lambda}$ are two lifts that define $\langle f\rangle$-equivariant stuctures on $\Lambda$. Thus $r=f_{\Lambda}^{-1}\circ f_{1}\in\operatorname{Aut}(\Lambda)\simeq\mathbbm{C}^{\times}$ must satisfy $r^{p}=1$. Since $f_{E}\mapsto\alpha\cdot f_{E}$ has the effect that $r\mapsto\alpha^{2}\cdot r$, we see that there is a unique normalization of $f_{E}$ (still defining an equivariant structure) such that $r=1$, which is equivalent to $h$ being an $\langle f\rangle$-equivariant isomorphism.
@@
Recall the notation $(\mathbb{E},\mathbb{\Phi})$ for the $\mathbb{T}$-equivariant universal Higgs bundle. The construction of the $\mathbb{T}$-equivariant structure is presented in [39, §3] and [37, §6.3.2]. The universal bundle is well-defined up to tensoring with a $\mathbb{T}$-equivariant line bundle pulled back from $\mathcal{M}$. Set $\mathbbm{U}=\operatorname{End}_{0}(\mathbb{E})$. We have that
@@
We consider ×$\mathcal{M}\times X$ a $\mathbb{T}\times\langle f\rangle$-variety, with the $\langle f\rangle$-action induced by the automorphism $F$ given by $F(\mathcal{E},x)=(f^{*}(\mathcal{E}),f^{-1}(x))$ for all $(\mathcal{E},x)\in\mathcal{M}\times X$. Consider the natural projection onto the second factor $\pi_{X}:\mathcal{M}\times X\rightarrow X$.
@@
There is a unique $\mathbb{T}\times\langle f\rangle$-equivariant structure on $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ such that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$. Let $f_{\operatorname{\mathbbm{E}}}$ denote the generator of the $\langle f\rangle$-action on $\operatorname{\mathbbm{E}}$ covering the action of $F$ on ×$\mathcal{M}\times X$. Through the isomorphism (20), the generator $f_{\operatorname{\mathbbm{E}}}$ induces the standard $\mathbb{T}\times\langle f\rangle$-equivariant structure on $T\mathcal{M}$.
@@
We will argue that $F_{\operatorname{\mathbbm{E}}}$ can be normalized to a new isomorphism which induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. Clearly $F^{p}=\operatorname{Id}$. It follows that $F_{\operatorname{\mathbbm{E}}}^{p}\in\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$. It is well-known that $\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ is one-dimensional. It follows that $F_{\operatorname{\mathbbm{E}}}^{P}=\lambda\cdot\operatorname{Id}$ for some $\lambda\in\mathbbm{C}\setminus\{0\}$. Thus we can replace $F_{\operatorname{\mathbbm{E}}}$ with $f_{\operatorname{\mathbbm{E}}}:=\lambda^{1/p}F_{\operatorname{\mathbbm{E}}}$ to obtain a lift that induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. Since $\gcd(p,2)=1$, the lift can be fixed by demanding that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$.
@@ 2.1 The Target Moduli Space of Parabolic Higgs Bundles
Let $[\cdot]_{p}:\mathbb{Z}\rightarrow\{0,...,p-1\}$ denote the remainder modulo $p$. For each $x\in X^{f}$, define $m_{x}\in\{1,...,p-1\}$ by the condition that $m_{x}n_{x}=-1\mod p$, where $n_{x}$ was defined in the introduction by the condition that $df_{x}=\zeta^{n_{x}}\cdot\operatorname{Id}$.
For each $x\in X^{f}$, let $b_{x}\in\{0,...,p-1\}$ be the unique solution to $\zeta^{2b_{x}n_{x}}=\alpha_{x}^{-1}$. Define the $\langle f\rangle$-invariant divisor $B=\sum_{x\in X^{f}}b_{x}$, and let $\mathcal{O}_{B}$ be the $\langle f\rangle$-equivariant line bundle associated with $B$, with the standard $\langle f\rangle$-equivariant structure as presented in Proposition 2.3. Then the eigenvalues of the $f$-action on $\Lambda(2B)$ over $X^{f}$ are all equal to unity. It follows that there exists a unique holomorphic line bundle of odd degree $\tilde{\Lambda}\in\operatorname{Pic}(\tilde{X})$ such that $\Lambda(2B)$ is $\langle f\rangle$-equivariantly isomorphic to $\pi^{*}(\Lambda)$ with the canonical $\langle f\rangle$-equivariant structure on the pullback line bundle. Namely, we can define $\tilde{\Lambda}$ as the quotient $\Lambda(2B)/\langle f\rangle$.
@@
  • Denote by ~$\tilde{\mathcal{M}}_{i}$ (resp. $\tilde{\mathcal{N}}_{i}$) the moduli space [46] of rank two stable parabolic Higgs bundles (resp. vector bundles) on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i},w_{i})$ with fixed determinant $\tilde{\Lambda}_{i}=\tilde{\Lambda}(D_{i}).$ Define ~$\tilde{\mathcal{M}}=\bigsqcup_{i\in\mathcal{I}}\tilde{\mathcal{M}}_{i}.$
For each $i\in\mathcal{I}$, the moduli space ~$\tilde{\mathcal{M}}_{i}$ is constructed as a hyper-Kähler quotient in [46] and proven to be a quasi-projective $\mathbb{T}$-variety in [69, 70] by means of GIT [50].
@@ 2.2 Hecke Transformations and Descent
Let $L$ be an $\langle f\rangle$-equivariant line bundle on $X$. Let $(E,F,\Phi)$ be an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle on $X$ such that $\det(E)$ is $\langle f\rangle$-equivariantly isomorphic to $L$. Fix $x\in X^{f}$. Assume that $\Phi$ is holomorphic in a neighbourhood of $x$. Assume that the eigenvalues of $f_{E}(x)$ are $1$ and $\zeta^{b_{x}}$, where $b_{x}\in\{1,...,p-1\}$, and assume $F_{x}$ is the the $\zeta^{b_{x}}$-eigenspace. Consider the $\langle f\rangle$-equivariant line bundle $L^{\prime}=L(-x)$ The following holds. There exists an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E^{\prime},F^{\prime},\Phi^{\prime})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $L^{\prime}$, together with a morphism of $\langle f\rangle$-equivariant quasi-parabolic Higgs bundles $\iota:(E^{\prime},F^{\prime},\Phi^{\prime})\rightarrow(E,F,\Phi)$ with the following properties.
@@
[…]
@@
The following facts are already covered in detail in [5]. The sequence (23) defines an $\langle f\rangle$-equivariant quasi parabolic bundle $(E^{\prime},F^{\prime})$ together with a vector bundle morphism $\iota:E^{\prime}\rightarrow E$, which is an isomorphism of $\langle f\rangle$-equivariant quasi parabolic bundles over $X\setminus\{x\}$. Further, one has an $\langle f\rangle$-equivariant isomorphism $\det(E^{\prime})\simeq\det(E)(-x)$. Thus, it remains to construct the Higgs field $\Phi^{\prime}$ and to prove the assertions about it. The Higgs field $\Phi^{\prime}$ is constructed on $X\setminus\{x\}$ by conjugating $\Phi$ by $\iota$. Thus it is clear that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle above $X\setminus\{x\}$, and it remains only to analyze $\Phi^{\prime}$ near $x$. This is done in two steps. First, we show that $\Phi^{\prime}$ extends meromorphically to $X$ with a potential pole of order one at $x$, and with residue being nilpotent with respect to the flag $F^{\prime}_{x}\subset E^{\prime}_{x}$. It immediately follows that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi parabolic Higgs bundle, because $\Phi^{\prime}$ is fixed by the $\langle f\rangle$-action on a dense subspace. In the second step, we verify the implication (22).
@@
Let $z$ be a local coordinate for $X$ centered at $x$. Choose a local frame $s=(s_{1},s_{2})$ of $E$ near $x$ such that with respect to this frame we have that $f_{E}(0)=\operatorname{diag}(1,\zeta_{b_{x}})$. Define $s^{\prime}=(s_{1}^{\prime},s_{2}^{\prime})=(s_{1},z\cdot s_{2})$. We now consider $\Phi^{\prime}$ near $x$. Let $(s_{1}^{*},s_{2}^{*})$ be the induced frame of $E^{*}$ near $x$, such that $s_{j}^{*}(s_{i})=\delta_{i,j}$ for $i,j\in\{1,2\}$, where $\delta_{i,j}$ is the Kronecker delta. For $i,j=1,2$ let $\phi_{i,j}\in\mathcal{O}_{X,x}$ be the regular function such that $\phi_{i,j}dz=s_{i}^{*}(\Phi(s_{j}))$. With respect to the frames $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ we see that $\iota$ is given by $\operatorname{diag}(1,z)$. With this notation, we have the following matrix decomposition of $\Phi^{\prime}$ with respect to the ordered frame $(s_{1}^{\prime},s_{2}^{\prime})$
@@
$\begin{split}f_{E^{\prime}}&=\begin{pmatrix}F_{1,1}&F_{1,2}\cdot z\\ F_{2,1}\cdot z^{-1}\zeta^{-n_{x}}&F_{2,2}\cdot\zeta^{-n_{x}}\end{pmatrix}.\end{split}$ (30)
@@
For a line subbundle $H\subset E$ we define $\Psi_{x}(H)$ to be equal to $H(-x)$ if $H_{x}\not=\operatorname{Im}(\iota(x))$, otherwise we define $\Psi_{x}(H)$ to be equal to $H$.
For an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}\rightarrow X$ on which $\langle f\rangle$ act as the identity above $X^{f}$, one can define a quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}/\langle f\rangle\rightarrow\tilde{X}$ by taking the quotient. We say $\operatorname{\mathcal{E}}/\langle f\rangle$ is obtained by descent. This is discussed in detail in the proof of Theorem 1.1 below.
For all $i\in\mathcal{I}_{0}$ define $\mathcal{M}_{i}\subset\mathcal{M}^{f}$ to be the $\mathbb{T}$-subvariety of $\mathcal{M}^{f}$ characterized by the fact that the eigenvalues of the $f$-action on $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i}\times X^{f}}$ are given by $i.$ It will follow from the proof of Theorem 1.2 that $\mathcal{M}_{i}$ is non-empty if and only if ~$\tilde{\mathcal{M}}_{i}$ is non-empty, i.e. if and only if $i\in\mathcal{I}$.
@@
Define the $\mathbb{T}$-equivariant morphism $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$ by
$\varXi_{i}(\operatorname{\mathcal{E}})=\left(\prod\limits_{x\in\mathcal{D}_{i}}\Psi_{x}^{[2m_{x}i_{x}]_{p}}(\operatorname{\mathcal{E}}(S_{i}))\right)/\langle f\rangle.$ (32)
It is easily seen that $\varXi_{i}(\mathcal{M}_{i}\cap\mathcal{N})=\tilde{\mathcal{N}}_{i}$, in accordance with the identification of the $\langle f\rangle$-fixed locus of the moduli space $\mathcal{N}$ of stable bundles given in [5], though we have used a slightly different isomorphism here.
Proof of Theorem 1.1.

### v2

@@ 1 The Fixed Locus
Let $Y$ be a compact Riemann surface with a finite subset $\delta$ of marked points. Our conventions and notation on quasi-parabolic Higgs bundles and parabolic Higgs bundles on $(Y,\delta)$ agree with those in [54], except we only consider full flags and specialize to the rank two case. Notice that the real numbers $\{\alpha_{j}(y)\}_{j=1,2,y\in\delta}$ introduced in Section $1$ in [54] satisfy that $\alpha_{j}(y)=w_{j}(y),j=1,2$ for all marked points $y\in\delta$.
@@
[…]
@@
where $f^{c}_{E}:\operatorname{End}_{0}(E)\rightarrow\operatorname{End}_{0}(f^{*}(E))$ is the isomorphism given by conjugation with $f_{E}$ and where $\tilde{f}^{*}(\Phi)$ is defined by equation (1). Notice that equation (2) is invariant under $f_{E}\mapsto\alpha\cdot f_{E}$ for any $\alpha\in\mathbbm{C}^{\times}$. Recall that since $(E,\Phi)$ is stable we have $f_{E}^{p}\in\operatorname{Aut}(E,\Phi)\simeq\mathbbm{C}^{\times}$. Thus, we can normalize to ensure $f_{E}^{p}=\operatorname{Id}$ and because of conjugation invariance, the equation (2) remain true. Thus $f^{*}([E,\Phi])=[(E,\Phi)]$ implies that there exists a lift $f_{E}$ of $f$ to $(E,\Phi)$. This is unique up to multiplying by an element of $\mu_{p}$. Now let $h:\det(E)\rightarrow\Lambda$ be an isomorphims (this exists by assumption). Let $f_{\Lambda}$ be the canonical $\langle f\rangle$-equivariant structure on $\Lambda$. Then $f_{1}:=h\circ\det(f_{E})\circ h^{-1}$ and $f_{\Lambda}$ are two lifts that define $\langle f\rangle$-equivariant stuctures on $\Lambda$. Thus $r=f_{\Lambda}^{-1}\circ f_{1}\in\operatorname{Aut}(\Lambda)\simeq\mathbbm{C}^{\times}$ must satisfy $r^{p}=1$. Since $f_{E}\mapsto\alpha\cdot f_{E}$ has the effect that $r\mapsto\alpha^{2}\cdot r$ and since $\gcd(2,p)=1$ we see that there is a unique normalization of $f_{E}$ (still defining an equivariant structure) such that $r=1$, which is equivalent to $h$ being an $\langle f\rangle$-equivariant isomorphism.
@@
Recall the notation $(\mathbb{E},\mathbb{\Phi})$ for the $\mathbb{T}$-equivariant universal Higgs bundle. The construction of the $\mathbb{T}$-equivariant structure is presented in [46, §3] and [43, §6.3.2]. The universal bundle is well-defined up to tensoring with a $\mathbb{T}$-equivariant line bundle pulled back from $\mathcal{M}$. Set $\mathbbm{U}=\operatorname{End}_{0}(\mathbb{E})$. We have that
@@
We consider $\mathcal{M}\times X$ a $\mathbb{T}\times\langle f\rangle$-variety, with the $\langle f\rangle$-action induced by the automorphism $F$ given by $F(\mathcal{E},x)=(f^{*}(\mathcal{E}),f^{-1}(x))$ for all $(\mathcal{E},x)\in\mathcal{M}\times X$. Consider the natural projection onto the second factor $\pi_{X}:\mathcal{M}\times X\rightarrow X$.
@@
There is a unique $\mathbb{T}\times\langle f\rangle$-equivariant structure on $(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ such that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$. Let $f_{\operatorname{\mathbbm{E}}}$ denote the generator of the $\langle f\rangle$-action on $\operatorname{\mathbbm{E}}$ covering the action of $F$ on $\mathcal{M}\times X$. Through the isomorphism (3), the generator $f_{\operatorname{\mathbbm{E}}}$ induces the standard $\mathbb{T}\times\langle f\rangle$-equivariant structure on $T\mathcal{M}$.
@@
We will argue that $F_{\operatorname{\mathbbm{E}}}$ can be normalized to a new isomorphism which induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. Clearly $F^{p}=\operatorname{Id}$. It follows that $F_{\operatorname{\mathbbm{E}}}^{p}\in\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$. It is well-known that $\operatorname{Aut}_{\mathbb{T}}(\operatorname{\mathbbm{E}},\mathbb{\Phi})$ is one-dimensional. It follows that $F_{\operatorname{\mathbbm{E}}}^{P}=\lambda\cdot\operatorname{Id}$ for some $\lambda\in\mathbbm{C}^{\times}$. Thus we can replace $F_{\operatorname{\mathbbm{E}}}$ with $f_{\operatorname{\mathbbm{E}}}:=\lambda^{1/p}F_{\operatorname{\mathbbm{E}}}$ to obtain a lift that induces a $\mathbb{T}\times\langle f\rangle$-equivariant structure. The lift can be fixed by demanding that $\det(\operatorname{\mathbbm{E}})$ is $\langle f\rangle$-equivariantly isomorphic to $\pi_{X}^{*}(\Lambda)$.
@@
Remark 2.
The $f$-action $f_{\operatorname{\mathbbm{E}}}$ on $\operatorname{\mathbbm{E}}$ from Proposition 2.2 covers $(f^{*},f^{-1}):\mathcal{M}\times X\rightarrow\mathcal{M}\times X$. Therefore, for every $E\in\mathcal{M}^{f}$ , the restriction of $f_{\operatorname{\mathbbm{E}}}$ to $\{E\}\times X$ is a lift $E\rightarrow(f^{-1})^{*}(E)$ that covers $f^{-1}:X\rightarrow X$. It follows that if the eigenvalues of the restriction of $f_{\operatorname{\mathbbm{E}}}$ to $\{E\}\times X^{f}$ are given by $x\mapsto[(\sqrt{\alpha_{x}}^{-1}\zeta^{i_{x}},\sqrt{\alpha_{x}}^{-1}\zeta^{-i_{x}})]$ for some $i\in\mathcal{I}_{0}$, then the eigenvalues of the corresponding restriction of the lift $E\rightarrow f^{*}(E)$ guaranteed by Proposition 1.1 are given by $x\mapsto[(\sqrt{\alpha_{x}}\zeta^{-i_{x}},\sqrt{\alpha_{x}}\zeta^{i_{x}})]$, and for each $x\in X^{f}$ and each eigenvalue $\eta$, the $\eta$-eigenspace of $f_{\operatorname{\mathbbm{E}}}$ restricted to $E_{x}$ is the $\eta^{-1}$-eigenspace of $f_{E}$ restricted to $E_{x}$.
@@ 1.1 The Target Moduli Space of Parabolic Higgs Bundles
Let $\cdot]_{p}:\mathbb{Z}\rightarrow\{0,...,p-1\}$ denote the remainder modulo $p$. For each $x\in X^{f}$, define $m_{x}\in\{1,...,p-1\}$ by the condition that $m_{x}n_{x}=-1\mod p$, where $n_{x}$ was defined in the introduction by the condition that $df_{x}=\zeta^{n_{x}}\cdot\operatorname{Id}$.
[…]
@@
  • Denote by $\tilde{\mathcal{M}}_{i}$ (resp. $\tilde{\mathcal{N}}_{i}$) the moduli space [54] of rank two stable parabolic Higgs bundles (resp. vector bundles) on $(\tilde{X},\operatorname{\tilde{\mathcal{D}}}_{i},w_{i})$ with fixed determinant $\tilde{\Lambda}_{i}=\tilde{\Lambda}(D_{i}).$ Define $\tilde{\mathcal{M}}=\bigsqcup_{i\in\mathcal{I}}\tilde{\mathcal{M}}_{i}.$
For each $i\in\mathcal{I}$, the moduli space $\tilde{\mathcal{M}}_{i}$ is constructed as a hyper-Kähler quotient in [54] and proven to be a quasi-projective $\mathbb{T}$-variety in [80, 81] by means of GIT [60].
@@ 1.2 Hecke Transformations and Descent
Let $L$ be an $\langle f\rangle$-equivariant line bundle on $X$. Let $(E,F,\Phi)$ be an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle on $X$ such that $\det(E)$ is $\langle f\rangle$-equivariantly isomorphic to $L$. Fix $x\in X^{f}$. Assume that $\Phi$ is holomorphic in a neighbourhood of $x$. Assume that the eigenvalues of $f_{E}(x)$ are $1$ and $\zeta^{b_{x}}$, where $b_{x}\in\{1,...,p-1\}$, and assume $F_{x}$ is the the $\zeta^{b_{x}}$-eigenspace. Consider the $\langle f\rangle$-equivariant line bundle $L^{\prime}=L(-x)$. The following holds. There exists an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $(E^{\prime},F^{\prime},\Phi^{\prime})$ with determinant $\langle f\rangle$-equivariantly isomorphic to $L^{\prime}$, together with a morphism of $\langle f\rangle$-equivariant quasi-parabolic Higgs bundles $\iota:(E^{\prime},F^{\prime},\Phi^{\prime})\rightarrow(E,F,\Phi)$ with the following properties.
@@
$0\rightarrow\mathcal{O}_{E^{\prime}}\overset{\iota}{\rightarrow}\mathcal{O}_{E}\overset{\lambda}{\rightarrow}T_{x}\rightarrow 0.$ (6)
@@
The following facts are already covered in detail in [5]. The sequence (6) defines an $\langle f\rangle$-equivariant quasi parabolic bundle $(E^{\prime},F^{\prime})$ together with a vector bundle morphism $\iota:E^{\prime}\rightarrow E$, which is an isomorphism of $\langle f\rangle$-equivariant quasi parabolic bundles over $X\setminus\{x\}$. Further, one has an $\langle f\rangle$-equivariant isomorphism $\det(E^{\prime})\simeq\det(E)(-x)$.
Thus, it remains to construct the Higgs field $\Phi^{\prime}$ and to prove the assertions about it. The Higgs field $\Phi^{\prime}$ is constructed on $X\setminus\{x\}$ by conjugating $\Phi$ by $\iota$. Thus it is clear that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle above $X\setminus\{x\}$, and it remains only to analyze $\Phi^{\prime}$ near $x$. This is done in two steps. First, we show that $\Phi^{\prime}$ extends meromorphically to $X$ with a potential pole of order one at $x$, and with residue being nilpotent with respect to the flag $F^{\prime}_{x}\subset E^{\prime}_{x}$. It immediately follows that $(E^{\prime},F^{\prime},\Phi^{\prime})$ is an $\langle f\rangle$-equivariant quasi parabolic Higgs bundle, because $\Phi^{\prime}$ is fixed by the $\langle f\rangle$-action on a dense subspace. In the second step, we verify the implication (LABEL:implication).
@@
Let $z$ be a local coordinate for $X$ centered at $x$. Choose a local frame $s=(s_{1},s_{2})$ of $E$ near $x$ such that with respect to this frame we have that $f_{E}(0)=\operatorname{diag}(1,\zeta_{b_{x}})$. The existence of such a frame is proven by elementary means in [5, Section 2]. Define $s^{\prime}=(s_{1}^{\prime},s_{2}^{\prime})=(s_{1},z\cdot s_{2})$. We now consider $\Phi^{\prime}$ near $x$. Let $(s_{1}^{*},s_{2}^{*})$ be the induced frame of $E^{*}$ near $x$, such that $s_{j}^{*}(s_{i})=\delta_{i,j}$ for $i,j\in\{1,2\}$, where $\delta_{i,j}$ is the Kronecker delta. For $i,j=1,2$ let $\phi_{i,j}\in\mathcal{O}_{X,x}$ be the regular function such that $\phi_{i,j}dz=s_{i}^{*}(\Phi(s_{j}))$. With respect to the frames $(s_{1},s_{2})$ and $(s_{1}^{\prime},s_{2}^{\prime})$ we see that $\iota$ is given by $\operatorname{diag}(1,z)$. With this notation, we have the following matrix decomposition of $\Phi^{\prime}$ with respect to the ordered frame $(s_{1}^{\prime},s_{2}^{\prime})$
@@
[…]
@@
For a line subbundle $H\subset E$ we define $\Psi_{x}(H)$ to be equal to H(x)H(-x) if $H_{x}\not=\operatorname{Im}(\iota(x))$, otherwise we define $\Psi_{x}(H)$ to be equal to $H$.
For an $\langle f\rangle$-equivariant quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}\rightarrow X$ on which $\langle f\rangle$ act as the identity above $X^{f}$, one can define a quasi-parabolic Higgs bundle $\operatorname{\mathcal{E}}/\langle f\rangle\rightarrow\tilde{X}$ by taking the quotient. We say $\operatorname{\mathcal{E}}/\langle f\rangle$ is obtained by descent. This is discussed in detail in the proof of Theorem LABEL:thm:M^f below.
For all $i\in\mathcal{I}_{0}$ define $\mathcal{M}_{i}\subset\mathcal{M}^{f}$ to be the $\mathbb{T}$-subvariety of $\mathcal{M}^{f}$ characterized by the fact that the eigenvalues of the $f$-action on $\operatorname{\mathbbm{E}}_{\mid\mathcal{M}_{i}\times X^{f}}$ are given by $i.$ It will follow from the proof of Theorem LABEL:ThmMain that $\mathcal{M}_{i}$ is non-empty if and only if $\tilde{\mathcal{M}}_{i}$ is non-empty, i.e. if and only if $i\in\mathcal{I}$.
@@
For all $x\in X^{f}$, define the positive integer $b_{i}(x):=[2m_{x}i_{x}]_{p}.$ Define the $\mathbb{T}$-equivariant morphism $\varXi_{i}:\mathcal{M}_{i}\rightarrow\tilde{\mathcal{M}}_{i}$ by
$\varXi_{i}(\operatorname{\mathcal{E}})=\left(\prod\limits_{x\in\mathcal{D}_{i}}\Psi_{x}^{b_{i}(x)}(\operatorname{\mathcal{E}}(S_{i}))\right)/\langle f\rangle.$ (13)
It is easily seen that $\varXi_{i}(\mathcal{M}_{i}\cap\mathcal{N})=\tilde{\mathcal{N}}_{i}$, in accordance with the identification of the $\langle f\rangle$-fixed locus of the moduli space $\mathcal{N}$ of stable bundles given in [5], though we have used a slightly different isomorphism here.We remark that a similar analysis involving the effect of Hecke transformations was considered by Witten in [77], though not in the $\langle f\rangle$-equivariant setting.
Proof of Theorem LABEL:thm:M^f.